OpenStudy (anonymous):

find two consecutive even intergers such that the larger, added to eight times the smaller, equals 110

OpenStudy (anonymous):

x,x+1 are your integers. use this to solve it

OpenStudy (anonymous):

, let x be the first integer, and x+1 the second

OpenStudy (anonymous):

so (x+1)+8x=110, can you take it from there?

OpenStudy (anonymous):

x+1+8x=110 9x=109 x=109/9... I think

OpenStudy (anonymous):

can u show me step bu step @fall2012

OpenStudy (anonymous):

weird... that x is not an integer

OpenStudy (anonymous):

sorry.. haha. x and x+2 are u nbers

OpenStudy (anonymous):

x+2+8x=110 9x=108 x=12

OpenStudy (anonymous):

were do u get the 9 from

OpenStudy (anonymous):

x+8x=9x

OpenStudy (anonymous):

sure. since the integers are consecutive even integers (thanks for picking that out @josiahh), it means that the larger integer would be the smaller integer plus two

OpenStudy (anonymous):

so your answer should be 12 and 14... thx @fall2012

OpenStudy (anonymous):

how did u get 108

OpenStudy (anonymous):

110-2=108

OpenStudy (anonymous):

but theres no - sign in the equation

OpenStudy (anonymous):

since the 8 times the larger integer (x+2) is equal to 110, we have (x+2)+8x=110

OpenStudy (anonymous):

yea but you have to subtract 2 from both sides

OpenStudy (anonymous):

so its 12 nd 14

OpenStudy (anonymous):

expanding the bracket, we have x+2+8x=110, then we collect like terms, so we have x+8x=110-2 9x=108, x=108/9 x=12 and the second integer will be 12+2,=14

OpenStudy (anonymous):

thanks u explained it better to my capacities